You are planning to save for retirement over the next 30 years. To save for retirement, you will invest $1,600 per month in a stock account in real dollars and $585 per month in a bond account in real dollars. The effective annual return of the stock account is expected to be 11 percent, and the bond account will earn 7 percent. When you retire, you will combine your money into an account with an effective return of 9 percent. The returns are stated in nominal terms. The inflation rate over this period is expected to be 3 percent.

a.

How much can you withdraw each month from your account in real terms assuming a 25-year withdrawal period? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

b. What is the nominal dollar amount of your last withdrawal? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Stock a/c balance after 30 years Px[(1+r)^n-1)=r Here, 1 Interest rate per annum 2 Number of years 3 Number of payments per p

Monthly withdrawal PVA:([1-(1=(1+r)^n)]=r) Here, A Interest rate per annum B Number of years C Number of compoundings per per

*Please rate thumbs up